OpenStudy (anonymous):

The number of trout t years later is given by P(t) = 700e^ − 0.4t when t=_______ years there are only 70 trout.

zepdrix (zepdrix):

So we have population of trout, `P`, as a function of time, `t`, in years.

zepdrix (zepdrix):

They want to know what value of t gives us: \(\Large\bf\sf P(t)=70\)

zepdrix (zepdrix):

\[\Large\bf\sf P(t)=700e^{-0.4t}\qquad\to\qquad 70=700e^{-0.4t}\]Understand that step? We're letting the population equal 70, and we want to figure out which t makes that happen.

zepdrix (zepdrix):

Oh and \(\Large\bf \color{#CC0033 }{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

yes I do.. i got to that step

OpenStudy (anonymous):

thank you :)

zepdrix (zepdrix):

So now we need to solve for t. It will involve using the natural long.

OpenStudy (anonymous):

okay

zepdrix (zepdrix):

Let's isolate the exponential first, we'll start by dividing each side by 700,\[\Large\bf\sf \frac{70}{700}=e^{-0.4t}\qquad\quad\to\qquad\quad \frac{1}{10}=e^{-0.4t}\]

zepdrix (zepdrix):

Then we'll take the log of each side,\[\Large\bf\sf \ln \frac{1}{10}=\ln e^{\color{orangered}{-0.4t}}\]Ok our rules of logs let's us do something with \(\Large\bf \color{orangered}{\text{this part}}\). Do you remember what it is?

OpenStudy (anonymous):

i don't remember...maybe it allow us to cancel e?

zepdrix (zepdrix):

Mmmm ya that's another way to look at it. But the rule I was thinking of was this:\[\Large\bf\sf \color{royalblue}{\log(a^b)\quad=\quad b\cdot \log(a)}\]Allows us to bring the exponent out front as a coefficient multiplier.

zepdrix (zepdrix):

\[\Large\bf\sf \ln \frac{1}{10}=\color{orangered}{-0.4t}\cdot\ln e\]

zepdrix (zepdrix):

And then ln e is simply 1.\[\Large\bf\sf \ln \frac{1}{10}=\color{orangered}{-0.4t}\]

OpenStudy (anonymous):

how did you get 1 for e?

zepdrix (zepdrix):

Not for e, for ln(e). Whenever you have a log in which the `base` matches the `contents`, the result is 1.\[\Large\bf\sf \ln (e)\quad=\quad \log_{\color{royalblue}{e}}(\color{royalblue}{e})\quad=\quad 1\]

OpenStudy (anonymous):

ohhhhh okay i got it :)

zepdrix (zepdrix):

Understand how to solve for t? c: It'll require some calculator work.\[\Large\bf\sf \ln(0.1)=-0.4t\]

OpenStudy (anonymous):

i got -2.305 for In(0.1)

OpenStudy (anonymous):

-2.303

OpenStudy (anonymous):

then i divide each side by -.04?

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

5.75

OpenStudy (anonymous):

is that correct?

zepdrix (zepdrix):

MMmmmmm yah that looks right! :) So in 5.76 years the population of trout will have fallen to 70.

zepdrix (zepdrix):

You can check your work by plugging t=5.76 into your original equation, it should give you a value close to 70

OpenStudy (anonymous):

okay! THANKKKKK YOU SO MUCH!! YOUR BETTER THAN MY MATH PROFESSOR!

zepdrix (zepdrix):

Haha give him a little credit XD It's hard to give every student personal attention :3

OpenStudy (anonymous):

ture! thank you so much!